AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 17:16:48 2025
Author
*
:
Subject
*
:
> > 07.51 DAQ continues files 22Feb/R97 > > ASIC settings 2022Jan31-11-52-00 > slow comparator 0x64 > > EG&G Ortec 448 Pulser OFF > > All system wide checks OK > > Attachment 1 - ADC data item stats > > Attachment 2 - FEE64 temps OK > > Attachments 3 & 4, 8 & 9 - per FEE64 1.8.W spectra 20us FSR and 200us FSR > > Attachments 5 & 6 - per FEE64 1.8.L spectra > > Attachment 7 - per FEE64 stat spectra > > Leakage current remains 8.36 uA, 100V. > > Bias power OFF, SIP ion pump power back ON. NIM crate powered off. Motors pi and nnrpi1 still on for remote access. > Compressed air off, water off. Fee relay power off and power cable disconnected. Motors box off. Safety pin for motors reinserted. > > Checked cable harnesses at cryring. 2 good short sections found and left with boxes in cryring. 2 damaged long sections (bad pin and broken ceramic) taken to repair in > Edinburgh.
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown